### 12 Mar 2024 v1

@@ 1 Introduction
Researchers frequently use publicly available data in policy analysis, which is usually provided at the aggregate level due to individual privacy concerns. For example, statewide standardized exam results are reported in the form of school-wide or school district-wide pass rates, as opposed to pass result and demographic information (e.g. gender, race, family income) for each individual student. Election data is available as vote shares and voter demographics over voting districts, as opposed to the vote and demographics information for each individual voter. However, it is well known that aggregate variables can be related in ways that are different from the same variables at the individual level, [^1] Firebaugh (1978) notes that “aggregate” and “individual” are relative terms, as individual variables correspond to the individual unit of analysis and need not correspond to an individual person. In the econometrics literature, the terms “macro” and “micro” are also often used. a problem known as an ecological fallacy. [^2] This problem has also been referred to as the problem of aggregation (Stoker, 1984).
In this paper I consider the problem of identifying linear combinations of conditional mean outcomes, $\mathbb{E}[Y_{i}|X_{1i},\dots,X_{Li}]$, when individual-level outcome $Y_{i}$ is binary and the only data that is observed is the marginal distribution of each individual-level variable over many groups, which I call aggregate data. I develop a partial-identification methodology that constructs sharp bounds by solving an optimization problem that considers all underlying joint distributions of individual-level covariates that are consistent with the observed marginal distributions. I also derive closed-form bounds on each conditional mean outcome $\mathbb{E}[Y_{i}|X_{1i},\dots,X_{Li}]$. Finally, I show how further restrictions on the underlying data generating process, like shape restrictions, can be incorporated into the optimization problem to obtain sharp bounds under those restrictions. Since I do not observe individual-level joint distributions, I develop valid inference procedures on the identified set using marginal information only.
To demonstrate how informative these bounds can be, I apply this methodology to several different simulated aggregate data sets, calibrated to a Rhode Island standardized exam score data set used later in an empirical application. I find that bounds become narrower on specific conditional outcomes when the observed , and wider when the observed . Imposing monotonicity shape restrictions in the empirical application helps narrow bounds on test score gaps, but bounds remain wide. This suggests that obtaining useful individual-level results from aggregate data alone requires further, stronger assumptions that should be carefully justified by the researcher.
@@
The ecological fallacy has been studied across many fields over the past several decades, beginning with the seminal work of Robinson (1950), Duncan and Davis (1953), Theil (1954), and Shiveley (1974). These early methods demonstrated issues with interpreting parameters from regressions with aggregate data as individual-level effects, and provided simple bounds on individual-level effects with a binary outcome and single binary covariate. Since then, research has been conducted on consideration of the identification of parameters (Stoker, 1984, 1986), testing for aggregation bias (Lee et al., 1990), empirical illustrations of inconsistent results when using aggregate as opposed to individual-level data (Hsiao et al., 2005), and many other topics in the sciences and social sciences beyond economics.
@@
While the problem of aggregation has been widely acknowledged as a major issue by the literature, to the best of my knowledge much of the literature is concerned with point identification of individual-level parameters (King et al., 1999; Rosen et al., 2001). [^3] One exception is Jiang et al. (2020), who develop partial identification techniques in a setting with two binary covariates and linear individual-level relationships. When only aggregate information is available, knowledge about parameters at the individual level is limited, and point identification is often not achieved without further assumptions. Imposing more assumptions may allow for precise results, but such assumptions may be less plausible. For example, one popular method for analyzing aggregate data is the ecological inference method (King, 1997), used often in political science studies of elections (Burden and Kimball, 1998). This method relies on many assumptions, like imposing individual-level joint distributions and a lack of bias introduced by the aggregation, that often fail to hold in applied settings (Tam Cho, 1998; Cho and Gaines, 2004).
@@
Even without these strong assumptions, partial identification is still possible and may reveal useful insights about the size or range of magnitude of parameters of interest. Partial identification in various contexts has been widely studied in the econometrics literature, especially over the last thirty years. [^4] See Tamer (2010), Ho and Rosen (2015), Molinari (2020), and Kline and Tamer (2023) for detailed surveys of partial identification in economics. Relevant to this paper is partial identification with data combination, since when combining two different data sets data the joint distribution between the data are unobserved (Cross and Manski, 2002; Molinari and Peski, 2006; Ridder and Moffitt, 2007; Fan et al., 2014, 2016). Relative to this literature, I assume that the joint distribution between every variable is unobserved and provide valid inference methods.
@@ 2.1 Identified set
Let $(Y_{i},X_{1i},\dots,X_{Li},G_{i}),i=1,\dots,n$ be an i.i.d. sequence of random variables. [^5] I assume individual-level variables are i.i.d. for the sake of simplicity; this assumption can be relaxed and the consistency and inference results of Section 3 will still hold. Suppose outcome $Y_{i}$ is binary, with $Y_{i}\in\{0,1\}$ and suppose $G_{i}\in\{1,\dots,G\}$ denotes the group of individual $i$. Further assume covariates $X_{i}\equiv(X_{1i},\dots,X_{Li})^{\prime}$ is discrete with known finite support $\{x_{k}\}_{k=1}^{K}\subseteq\mathbb{R}^{L}$. [^6] The case with multinomial or continuous outcome and continuous covariates is beyond the scope of this paper.
The goal is to construct bounds on linear combinations of $\mathbb{E}[Y_{i}|X_{i}=x_{k}]$, $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$, for given weights $\{\lambda_{k}\}_{k=1}^{K}$. For example, if we are interested in the average marginal effect of changing $X_{i}$ from $x_{k_{1}}$ to $x_{k_{2}}$ on $Y_{i}$, we would choose $\lambda_{k_{2}}=1,\lambda_{k_{1}}=-1,$ and $\lambda_{k}=0$ for all other $k$. I will construct identified sets using only expressions for $\mathbb{E}[Y_{i}|G_{i}=g],\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g],$ and $\mathbb{P}[G_{i}=g]$ for every $\ell=1,\dots,L,k=1,\dots,K,g=1,\dots,G$; the sample equivalents of these parameters are observed in aggregate data.
@@
I summarize the previous statements in the following assumption:
@@
    $Y_{i}$ is binary.
@@
I defer all proofs to Appendix A. Proposition 1 states that we can equivalently express the identified set $D$ given by (4) as the weighted sum of solutions to bilevel optimization problems. This formulation is helpful because it suggests how computation of the lower and upper bound might be performed. In particular, solving for

### 05 Apr 2024 v3

@@ 1 Introduction
Researchers frequently use publicly available data in policy analysis, which is usually provided at the aggregate level due to individual privacy concerns. For example, statewide standardized exam results are reported in the form of school-wide or school district-wide pass rates, as opposed to pass result and demographic information (e.g. gender, race, family income) for each individual student. Election data is available as vote shares and voter demographics over voting districts, as opposed to the vote and demographics information for each individual voter.
Many researchers use aggregate data to run analyses where the outcome of interest is at the individual level. One example is the literature on the relationship between religion and suicide, which often uses suicide rates and religion prevalence rates (Neeleman et al., 1997; Dervic et al., 2004; Becker and Woessmann, 2018). Jack et al. (2023) looks at the relationship between virtual learning during the COVID-19 pandemic and student performance in grade school using standardized exam pass rates and district demographic information. Alter et al. (1999) looks at the relationship between median neighborhood income, rate of use of cardiac procedures, and mortality rates after a heart attack to determine the relationship between patient income and treatment outcomes. Examples of aggregate data used in political science analyses of individual voting behavior are given in Freedman et al. (1998) and Kousser (2001). Another large literature studies the relationship between mortality and economic conditions and often utilizes aggregate data in the form of mortality rates and unemployment rates (Ruhm, 2000; Neumayer, 2004; Gerdtham and Ruhm, 2006; Svensson, 2007; Lindo, 2015).
@@
It is well known that aggregate variables can be related in ways that are different from the same variables at the individual level, [^1] Firebaugh (1978) notes that “aggregate” and “individual” are relative terms, as individual variables correspond to the individual unit of analysis and need not correspond to an individual person. In the econometrics literature, the terms “macro” and “micro” are also often used. a problem known as an ecological fallacy. [^2] This problem has also been referred to as the problem of aggregation (Stoker, 1984). The ecological fallacy has been studied across many fields over the past several decades, beginning with the seminal work of Robinson (1950); Duncan and Davis (1953), and Theil (1954). These earlier papers demonstrated issues with interpreting parameters from regressions with aggregate data as individual-level effects, and provided simple bounds on individual-level effects with a binary outcome and single binary covariate. Since then, studies like Shiveley (1974); Kramer (1983); Greenland and Robins (1994); Gravelle et al. (2002); Hsiao et al. (2005), and Lindo (2015) have continued to demonstrate that results on individual-level effects are sensitive to the level of data aggregation and other imposed assumptions like functional form.
@@
While the ecological fallacy has been acknowledged as an issue by the literature, to the best of my knowledge much of the literature is concerned with point identification of individual-level parameters (King et al., 1999; Rosen et al., 2001). [^3] One exception is Jiang et al. (2020), who develop partial identification techniques in a setting with two binary covariates, imposing linear individual-level relationships. When only aggregate information is available, knowledge about parameters at the individual level is limited, and point identification is often not achieved without further assumptions. Imposing more assumptions may allow for precise results, but such assumptions may be less plausible. For example, one popular method for analyzing aggregate data is the ecological inference method (King, 1997), used often in political science studies of elections (Burden and Kimball, 1998). This method relies on many assumptions, like imposing individual-level joint distributions and a lack of bias introduced by the aggregation, that often fail to hold in applied settings (Tam Cho, 1998; Cho and Gaines, 2004; Freedman et al., 1998; Kousser, 2001).
Even without these strong assumptions, partial identification is still possible. The resulting identified set is exactly the extent to which individual-level results are sensitive to assumptions. In this paper I consider the problem of identifying linear combinations of conditional mean outcomes, $\mathbb{E}[Y_{i}|X_{1i},\dots,X_{Li}]$, when individual-level outcome $Y_{i}$ is binary and the only data that is observed is the marginal distribution of each individual-level variable over many groups, which I call aggregate data. I develop a partial identification methodology that constructs sharp bounds by solving an optimization problem that considers all underlying joint distributions of individual-level covariates that are consistent with the observed marginal distributions. I show how further restrictions on the underlying data generating process, like shape restrictions or additional data at a finer level of aggregation, can be incorporated into the optimization problem to obtain sharp bounds. Since I do not observe individual-level joint distributions, I develop valid inference procedures on the identified set using marginal information only.
To demonstrate how informative these bounds can be, I apply this methodology to several different simulated aggregate data sets calibrated to the Rhode Island standardized exam score data set used in my empirical application. I find that bounds are relatively wide on marginal effect parameters of interest. Imposing monotonicity shape restrictions in the empirical application helps narrow bounds on test score gaps, and using additional data at a finer level of aggregation makes bounds on test score gaps even narrower. This suggests that it is easier to obtain useful individual-level results from aggregate data when the data is aggregated at a finer level and when the researcher can impose more restrictive assumptions about the individual-level data generating process.
@@
Partial identification in various contexts has been widely studied in the econometrics literature, especially over the last thirty years. [^4] See Tamer (2010), Ho and Rosen (2015), Molinari (2020), and Kline and Tamer (2023) for detailed surveys of partial identification in economics. Relevant to this paper is partial identification with data combination, since when combining two different data sets the joint distribution between the data sets is often unobserved (Cross and Manski, 2002; Molinari and Peski, 2006; Ridder and Moffitt, 2007; Fan et al., 2014, 2016). However in the data combination literature the joint distribution within each data set is observed; in the aggregate data setting the joint distribution between every combination of variables is unobserved. In particular, the setting in Cross and Manski (2002) applies when we observe the marginal distribution of $Y_{i}$ over groups and the joint distribution of covariates $X_{i}=(X_{1i},\dots,X_{Li})$ over groups. Aggregate data often does not fit this setting; for example aggregate data can provide ethnicity distributions, gender distributions, education distributions, and income distributions but will not provide joint distributions of all of these variables, often due to privacy concerns. Being able to observe the joint distribution of covariates makes bounds narrower than when we do not observe the joint distribution, as further discussed in Appendix A.3. The intuition for this is that since individual-level effects usually depend on the joint distribution of all variables, being able to pin down the joint distribution of most of the variables will restrict the size of the identified set.
@@ 2.1 Identified set
Let $(Y_{i},X_{1i},\dots,X_{Li},G_{i}),i=1,\dots,n$ be a sequence of random variables. Suppose outcome $Y_{i}$ is binary, with $Y_{i}\in\{0,1\}$ and suppose $G_{i}\in\{1,\dots,G\}$ denotes the group of individual $i$. Further assume covariates $X_{i}\equiv(X_{1i},\dots,X_{Li})^{\prime}$ are discrete with known finite support $\{x_{k}\}_{k=1}^{K}\subseteq\mathbb{R}^{L}$. [^5] The case with multinomial or continuous outcome and continuous covariates is beyond the scope of this paper.
The goal is to construct bounds on linear combinations of $\mathbb{E}[Y_{i}|X_{i}=x_{k}]$, $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$, for given weights $\{\lambda_{k}\}_{k=1}^{K}$. For example, if we are interested in the average marginal effect of changing $X_{i}$ from $x_{k_{1}}$ to $x_{k_{2}}$ on $Y_{i}$, we can choose $\lambda_{k_{2}}=1,\lambda_{k_{1}}=-1,$ and $\lambda_{k}=0$ for all other $k$. I will construct identified sets using only expressions for $\mathbb{E}[Y_{i}|G_{i}=g],\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g],$ and $\mathbb{P}[G_{i}=g]$ for every $\ell=1,\dots,L,k=1,\dots,K,g=1,\dots,G$; the sample equivalents of these parameters are observed in aggregate data.
@@
In what follows I will maintain the following assumption:
@@
    $G_{i}$ is discrete with finite support $\{1,\dots,G\}$.
@@
    $G_{i}$ is i.i.d. and $(Y_{i},X_{i})|G_{i}$ are i.i.d. [^6] I assume individual-level variables are (conditionally) i.i.d. for the sake of simplicity; this assumption can be relaxed and the consistency and inference results of Section 3 will still hold.
@@
I defer all proofs to Appendix B. Proposition 1 states that we can equivalently express the identified set $D$ given by (4) as the weighted sum of solutions to bilevel optimization problems. This formulation is helpful because it suggests how computation of the lower and upper bound might be performed. In particular, solving for